# Experimental verification of the quantum nature of a neural network

### v4

@@ Abstract
Neural networks are being used to improve the probing of the state spaces of many particle systems as approximations to wavefunctions and in order to avoid the recurring sign problem of quantum monte-carlo. One may ask whether the usual classical neural networks have some actual hidden quantum properties that make them such suitable tools for a highly coupled quantum problem. I discuss here what makes a system quantum and to what extent we can interpret a neural network as having quantum remnants. I suggest that a system can be quantum both due to its fundamental quantum constituents and due to its dynamics, therefore, we can obtain entanglement both due to the quantum constituents’ nature and due to the functioning rules, or, in category theory terms, both due to the quantum nature of the objects of a category and of the maps. From a practical point of view, I suggest a possible experiment that could extract entanglement from the dynamics (maps) of an otherwise cvasi-classical (from the point of view of the constituents) neural network.
@@ I Introduction
The goal of this article is to show that the dynamics of a cvasi-classical neural network may have underlying quantum properties that can be amplified by its dynamics. It must be noted that the neural network referred to in this article is not a numerical simulation, as a simulation would ultimately, by design, have no complex phase and hence would not present quantum effects. My discussion is based on a real physically implemented cvasi-classical neural network. The main observation is that such a neural network may generate quantum effects due to the dynamics (or, in a categorical sense, due to the maps between the objects, not the objects themselves). The end result is that a cvasi-classical neural network, due to its dynamics, could generate entangled states and could be used for entanglement generation and extraction. As such, this is of particular interest to subjects like quantum biology or to describing quantum phenomena in natural neural networks where the usual criteria of temperature and scale would otherwise block such effects. Probably the best way of starting this article is to ask a question: what does it mean to be quantum? Lacking a proper axiomatisation of quantum mechanics this question is particularly important because systems or mathematical structures we may usually associate to classical physics may have unexpected quantum properties. From another point of view, systems usually modelled by means of strictly classical rules, may, upon physical implementation present quantum behaviour that escaped our previous analysis. Quantum mechanics has several defining principles, two of the most important being: its probabilistic nature: the outcomes generated by quantum mechanical calculations are always probabilistic, even if that probability gets very close to 0 or 1, and second, there is no pre-determined uniquely defined state of a system in the absence of an observable capable of asking a question about that property of the system. This property makes the probability amplitude and hence the wavefunction of the system represent the maximal knowledge attainable about the system in a given context. Essentially, that means the quantitative properties of a state of a system are generally not uniquely defined, and sometimes are not even defined at all in the absence of an observable to be associated to a method of determining them. Moreover, not any observable is capable of fully determining the numerical value of a property of the state of a system, and some observables may not be compatible with respect to a given property. As a result, some of the matrix valued observables of quantum mechanics may not commute with each other. Due to these properties, the wavefunction is, somehow unfortunately, called ”non-realistic”, and that is unfortunate because there is nothing ”unreal” about the wavefunction, in fact it provides the maximum of information about the system, and in many cases far more than a classical system could ever hope to achieve. A wavefunction ought to be interpreted as a catalogue of maximal knowledge that can be obtained about a property of the system to be analysed. As a consequence, the rules of obtaining the probabilistic outcomes of quantum mechanics change. We therefore have to employ a method of obtaining a module of a complex number, from a fundamentally complex quantity, the wavefunction. This wavefunction needs to be complex in order to account for correlations in our knowledge about the system. In turn, by this method we will generate the well known interference patterns in the probabilistic outcomes of quantum mechanics. That last property of quantum mechanics is Born’s rule.
@@
There exists a somehow widespread assumption that quantum phenomena must be either very small, or happen at low enough energy, or be small corrections of the order of $O(\hbar)$ to classical phenomena and that there is some form of ”gap” between a so called ”quantum world” and the ”real” or ”every-day” world. My thesis is that these assumptions are wrong. The approaches to quantum mechanics phenomena initiated up to now are based on controlling the interference of the ”catalogue of knowledge” wave patterns (basically, a sort of wavefunction engineering). Those methods are extremely susceptible to quantum decoherence, namely extended interaction and entanglement with the environment which leads to an exponential increase of the number of terms arising in the ”catalog of knowledge” describing the environment and the system. If one approaches quantum phenomena from a categorical perspective, one notices that the maps between quantum objects may as well have quantum properties, may become non-cartesian, and can produce and exploit entanglement. The maps between quantum objects are associated to the specific dynamics through which the system goes, and this very dynamics can produce, enhance, and utilise entanglement.
@@
One direction of research nowadays is to find some classical underpinnings for quantum mechanics. This research is an important part of the so called ”foundational” research on quantum mechanics. While the common approach is to lift or extend the observables of a system from the uniquely defined classical real numbers to matrix valued observables that encode several possible outcomes without the possibility of predicting which one will occur from the data about the system itself, some quantum foundational research tries to find a classical (therefore well determined and unique) way of characterising the system based on some form of hidden classical variables. My approach is the exact opposite of this path. Instead of searching for some underpinning classical structure to quantum mechanics, my view is that even systems that appear as classical, may still have hidden quantum properties, due, for example to the quantum nature of the transformation maps between the objects of the specific category considered. In particular, systems assumed to be classical, like a neural network, despite being simulated as classical objects in conventional computers, when physically realised, may exhibit quantum phenomena as cvasi-classical systems. In particular, the non-linear dynamics of neural networks may in fact amplify quantum behaviour.
@@
The assumption I wish to challenge is that quantum phenomena are strictly the result of objects being ”quantum” as opposed to ”classical”. Instead of a classical interaction between systems having well determined states, conventionally, one is expected to transition to a generalised interaction in which some classical object is replaced by a ”quantum” object for which not all of its properties are uniquely defined. This is the origin of the requirement of integrating over inner loop momenta in Feynman diagrams or the construction of the Feynman path integral. This is of course correct, but it is not the only way quantum phenomena may emerge. Indeed, I wish to show that quantum phenomena can emerge strictly from the rules used to describe the dynamics of cvasi-classical systems. Those rules themselves, seem more fundamental and can make the system generate quantum (or at least, almost quantum) behaviour. In this sense, even if the universe started with strictly classical objects and information (which is probably not the case), there exist certain rules of manipulating this classical information that will result in quantum behaviour. The quantum nature seems therefore embedded in the very rules of the system’s dynamics and does not need a separation between ”quantum” and ”classical” objects. The rules can be seen in categorical terms as maps, and the view drafted above would focus on the maps instead of the objects of a category. In the case of a cvasi-classical neural network, those rules would be those of the dynamical behaviour of the network.
@@
There are several macroscopic systems that are generally assumed to be classical, that by the way in which they work, have structures that give them some quantum properties, as far as the definitions above are concerned. One of these systems is the neural network. It is important to mention that there have been essays in which one speculated about the possibility of generating quantum computing processes in big enough biological neural networks by focusing on some detailed mechanism in some molecular substructure of their neural network [1]. I think such an approach is not feasible on a large scale (although one has discovered entanglement of spins in some animal brains [2,3] and substantial progress is being done in understanding macroscopic quantum phenomena from this point of view). Those methods are extremely susceptible to decoherence. However, one may ask what if the large structure of a network has properties that at least in part reflect the fundamental axioms of quantum mechanics? There have also been essays in constructing so called quantum neural networks [4]. Those networks have the links represented by quantum processes implemented by unitary transformations, turning the neural network into a large quantum circuit. Those methods are promising and it would be amazing if they will be achieved. However, there is still the problem of decoherence that will require such networks to be extremely robust, unfortunately a property still beyond today’s reach. It has however been noticed that classical neural networks do a great job at solving many body quantum problems with strong coupling [5], that are fundamentally quantum and expected to be properly solved on quantum computers. One role neural networks play in solving strongly correlated quantum systems is through simplifying the wavefunction by providing us with a more compact representation thereof. Another role played by neural networks is to improve the understanding of quantum dynamics [12-16]. What are the properties of a classical neural network that would make it suitable for solving such complex quantum problems? While it is true that classical neural networks seem to be very suitable for simulating quantum processes, I wish to ask in this article whether this is due to the fact that the rules of functioning of a neural network that appears classical from the perspective of its objects (namely neurons) has in fact hidden quantum properties at the level of the morphisms (maps) between those objects.
@@
Indeed, various quantum properties appear to be general properties not only of physical particles, but also of a series of mathematical structures that have a broad applicability. Non-separability, as encountered in the definition of entanglement, is also found, under the guise of non-cartesianity in the construction of fibre bundles, in which basically the bundle measures the global failure of maintaining the local cartesianity of a certain pairing. A non-trivial fibre bundle is defined to characterise precisely such a failure of the cartesian product between the fibre and the base to expand globally. This is just a mathematical example that could make the reasoning I try to follow here clearer: quantum properties are not bound to particles, but can also characterise various rules by which dynamical systems function. Such quantum properties make them function differently compared to their classical counterparts. Of course, this doesn’t mean a classical fibre bundle is quantum. We need additional structure for this. We need some type of quantum polarisation (a la geometric quantisation) and some form of half form correction, both general properties of quantisation, but there exists a deep similarity between the category of cobordisms whose objects are (n-1)-dimensional manifolds and its morphisms are n-dimensional cobordisms, and the catgeory of Hilb whose objects are Hilbert spaces used to describe states and the morphisms are linear operators describing ”processes” [32]. With this relation, the cartesian non-separability property that leads to entanglement in quantum mechanics has a similar role in other situations that involve non-separability, as is the case in fibre bundles. From a practical perspective, it is important to know that such entanglement exists in a category related by a functor to the initial one, as it would be possible to harness it as a quantum resource and use it in circumstances that are apparently remote from the initial categorical structure. In particular, an apparently classical neural network may extract its entanglement from an associated category and behave like a quantum system when its dynamics is analysed. A wavefunction, or a probability amplitude (a precursor to the probability we are used to in classical physics) can be associated not only to the objects of the category, but also to the maps between those objects, and hence superposition and even entanglement described at the level of the wavefunction associated to the quantum maps may prove to be a new quantum resource that could be harnessed by physical neural networks. This would lead to the possibility of having quantum computation like effects occurring even in the most noisy and hot environments, like the ones existing in animal brains. A physical feature emerging from the quantum properties above is the inseparability of state spaces of quantum states [8]. This obstruction to cartesian pairing has been traditionally called entanglement, and it is advisable to keep that notion even when applied to more abstract mathematical objects [9], [32].
@@
Entanglement is considered a characterising feature of quantum mechanics. It is a clearly non-classical phenomenon capable of offering correlations above any that could be obtained classically [10]. Entanglement is based on the non-separability of state spaces of systems that are connected to form an over-arching super-system. The non-separability relies on two concepts. First, taking two systems, each having their state spaces given, by combining them one obtains a paired state space with a dimension larger than the sum of the dimensions of the two subspaces, actually, the dimension will grow as the product of the dimensions of the two subspaces, leading to a pairing that involves significant global information that cannot be mapped or retrieved by local measurements on any of the two subsystems separately. The fact that pairing is not always cartesian is a feature of quantum mechanics that has correspondences in many algebraic structures. One can link this with the commutation relations that become non-abelian for incompatible observables, and finally with a probabilistic interpretation that allows, by means of statistics, to determine interference patterns linked to the global structure of the manifold on which the phenomena occur. In all cases, the two main aspects of the ”quantumness” of a system are the linear product structure, and its ability to provide access to non-local features by means of a probabilistic, wave-function based interpretation. It is worth noting that the entangled states are, as any quantum state, maximal states of knowledge. They are being defined to describe global properties of the combined system in the case in which separated independent descriptions of each of the subsystems that are combined are unachievable and don’t even exist for the specified observables in an unambiguous sense.
@@
In this article I show that a (cvasi)-classical neural network can indeed be represented by a non-cartesian monodromic category and that there exists a functor that links it to the $Hilb$ category. By doing so, a normal (cvasi)-classical neural network will show properties of quantum entanglement at a large scale in the same way in which other quantum systems show. However, being highly decoherent at the level of the constituent objects, it will be up to the maps to act as a resource for quantum entanglement. This means that the quantum nature emerges from the rules we impose on how the network should work, not from the actual quantum nature of its constituents, but finally, there should be no difference between a ”natural” quantum system made out of fundamental constituents that are ”quantum”, and one that is forced to function by rules (maps or morphisms of the category that are quantum) that result in an overall quantum behaviour of the network. The parameters of the network can be encoded in a Hilbert Kernel, where the respective parameters play the role of linear superpositions of the learned results, similar to the quantum superpositions allowed by the complex phases of a wavefunction. Therefore, there appears to be a strong analogy (if not more, a functorial connection) between entangled quantum states and neural networks and between the category of Hilbert spaces and the supposed category of neural networks.
From a categorical point of view, the general approach to quantum mechanics is that the quantum nature emerges from properties of the category-theory objects, which in physics would be electrons, photons, etc. My approach tries to show that while this is certainly true, we can as well, and in a sense in parallel, also have quantum phenomena emerging from the category-theoretical maps, linking objects. I must underline that this doesn’t invalidate the other approach. We can continue to talk equally well about the quantum nature of photons, electrons, or gravitons. However, in some cases, say, in the case of neural networks, as a simple example, we may have to think of the quantum nature emerging from category-theory maps instead of objects. The word ”emergence” may as well be used with caution, as in physics emergent properties are usually understood as properties emerging from the statistical dynamics of microscopic constituents. This is not how I used the term here. I rather refer to the emergence of quantum properties in a purely mathematical sense, as a result of properties of maps, instead of the properties of objects.
The distinction between a neural network as an ”unphysical” computational phenomenon as opposed to any other ”physical” phenomenon seems rather artificial. Any structure, mathematical or physical can be characterised as ”quantum” either because it has some intrinsic quantum nature (making some of the properties we expect it to have, undefined or not uniquely defined) or because the rules it must obey have to be such that the quantum nature emerges (becoming quantum due to the maps instead of objects). It is worth noting that recent research has succeeded in implementing a physical feed-forward neural network based on a series of physical objects that can be to some extent tuned to have only classical behaviour. Indeed it may be possible to experimentally test the morphisms origin of some quantum behaviour of a neural network by looking at a physical implementation of a neural network that appears to be classical from the perspective of the objects involved, but may gain quantum properties due to its inherent morphisms. Such physical neural networks have been constructed [17]. Until such an experiment can be performed, I can offer a numerical simulation showing that indeed this is the case, namely that provided with proper input, a physical, classical neural network can entangle the particles that it receives as input.

### v5

@@ Abstract
Neural networks are being used to improve the probing of the state spaces of many particle systems as approximations to wavefunctions and in order to avoid the recurring sign problem of quantum monte-carlo. One may ask whether the usual quasi-classical neural networks have some actual hidden quantum properties that make them such suitable tools for a highly coupled quantum problem. I discuss here what makes a system quantum and to what extent we can interpret a neural network as having quantum remnants. I suggest that a system can be quantum both due to its fundamental quantum constituents and due to its dynamics (seen as maps in a category), therefore, we can obtain entanglement both due to the quantum constituents’ nature and due to the functioning rules, or, in category theory terms, both due to the quantum nature of the objects of a category and of the maps. From a practical point of view, I suggest a possible experiment that could extract entanglement from the dynamics (maps) of an otherwise cvasi-classical (from the point of view of the constituents) neural network.
@@ I Introduction
The goal of this article is to show that a quasi-classical neural network may have underlying quantum properties that can be amplified by its dynamics. It must be noted that the neural network referred to in this article is not a numerical simulation, as a simulation would ultimately, by design, have no complex phase and hence would not present quantum effects. My discussion is based on a real physically implemented quasi-classical neural network. The main observation is that such a neural network may generate quantum effects due to the dynamics (or, in a categorical sense, due to the maps between the objects, not the objects themselves). The end result is that a quasi-classical neural network, due to its dynamics, could generate entangled (quantum) states and could be used for entanglement generation and extraction. As such, this is of particular interest to subjects like quantum biology or in describing quantum phenomena in natural neural networks where the usual criteria of temperature and scale would otherwise block such effects.
@@
Probably the best way of starting this article is to ask a question: what does it mean to be quantum? This question is particularly important because systems or mathematical structures that we may usually associate to quasi-classical physics may have unexpected quantum properties. From another point of view, systems usually simulated by means of strictly classical rules may, upon physical implementation, present quantum behaviour that escaped our previous analysis. Quantum mechanics has several defining principles, two of the most important being: its probabilistic nature: the outcomes generated by quantum mechanical calculations are always probabilistic, even if that probability gets very close to 0 or 1, and second, there is no pre-determined uniquely defined state of a system in the absence of an observable capable of asking a question about that property of the system. This property makes the probability amplitude and hence the wavefunction of the system represent the maximal knowledge attainable about the system in a given context. The wavefunction, being subject to a linear differential equation, allows the usage of the principle of superposition to find new solutions, with the consequence that physical systems, described by such probability amplitudes, can exist in what is known as superposed states. Moreover, the quantitative properties of a state of a system are generally not uniquely defined, and sometimes are not even defined at all in the absence of an observable to be associated to a method of determining them. Not any observable is capable of fully determining the numerical value of a property of the state of a system, and some observables may not be compatible with respect to a given property. As a result, some of the matrix valued observables of quantum mechanics may not commute with each other. Due to these properties, the wavefunction is, somehow unfortunately, called ”non-realistic”, and that is unfortunate because there is nothing ”unreal” about the wavefunction, in fact it provides the maximum of information about the system, and in many cases far more than a classical system could ever hope to achieve. A wavefunction ought to be interpreted as a catalogue of maximal knowledge that can be obtained about a property of the system to be analysed. As a consequence, the rules of obtaining the probabilistic outcomes of quantum mechanics change with respect to a classical system. We therefore have to employ a method of obtaining a module of a complex number, from what is a fundamentally complex quantity, namely the wavefunction. This wavefunction needs to be complex in order to account for correlations in our knowledge about the system. In turn, by this method we will generate the well known interference patterns in the probabilistic outcomes of quantum mechanics, by which quantum mechanics differs from classical mechanics. That last property of quantum mechanics is Born’s rule.
One of the main distinctions between classical and quantum mechanics is therefore that in classical mechanics probabilities are associated strictly to lack of knowledge about otherwise well expressed and physically realised properties of the system, while in quantum mechanics, probabilities emerge out of a lack of knowledge that is inherent to the system, and for which there is no physically realised property of the system to talk of.
The distinction between these two situations is rather formal, as, of course, nobody could in principle measure a system as being in a superposed state a-priori. To even realise the existence of such superposed states, one would have to perform a statistical experiment and see how the distributions differs from that of a system that we would imagine to satisfy the rules of classical probabilities. Indeed, such statistical experiments have been performed and they overwhelmingly prove the validity of quantum mechanics and the existence of a wavefunction based description of nature.
This however leaves open the possibility that systems that we imagined to be classical are in fact only quasi-classical and their quantum properties are somehow suppressed in the specific context we constructed for them. This would also rise the question of whether it would be possible by some means, to amplify such quantum properties in otherwise quasi-classical systems so that they can perform certain tasks (like, for example entangling some states) that could become relevant in quantum computers. This observation is not new. Indeed [37] shows that large structures regain their quantum properties and act as quantum objects in a double slit experiment, given a suitable parametric setup of the experiment.
There exists a somehow widespread assumption that quantum phenomena must be either very small, or happen at low enough energy, or be small corrections of the order of $O(\hbar)$ to classical phenomena and that there is some form of ”gap” between a so called ”quantum world” and the ”real” or ”every-day” world. Moreover, one usually assumes that the limit taken to obtain the classical limit of a quantum system involves trivially taking $\hbar\rightarrow 0$ as is commonly described using the Ehrenfest approximation to calculate expectation values. My thesis is that these assumptions are wrong. It is well known that, by Ehrenfest theorem, the relation to classical mechanics is trivially obtained only when the geometry of the system allows for a well localised wavefunction, which is not the case in most situations of experimental relevance in quantum mechanics. Moreover, the situation changes dramatically when the classical limit involves non-linear dynamics as is the case in quasi-classical physically realised neural networks. In those cases, the relation between quantum dynamics and their quasi-classical limit is in fact much more complicated and interesting.
@@
The approaches to quantum computing discussed up to now are based on controlling the interference of the ”catalogue of knowledge” wave patterns (basically, a sort of wavefunction engineering). Those methods are extremely susceptible to quantum decoherence, namely extended interaction and entanglement with the environment which leads to an exponential increase of the number of terms arising in the ”catalog of knowledge” describing the combined system of environment and computational device If one approaches quantum phenomena from a categorical perspective, one notices that the maps between quantum objects may as well have quantum properties, may become non-cartesian, and can produce and exploit entanglement. The maps between quantum objects are associated to the specific dynamics through which the system goes, and this very dynamics can produce, enhance, and utilise entanglement.
In particular the non-linear dynamics associated to neural networks appears to have an interesting effect on the underlying quantum properties. One may think that I am discussing two different types of dynamics in an equivalent way, namely on one side the classical dynamics of a neural network, where the non-linearities occur, and on the other side, a quantum dynamics of a quasi-classical neural network seen as some limit of a pure quantum situation that behaves linearly according to the laws of quantum mechanics. While this may appear so if one thinks of nature as being strictly divided between a classical and a quantum realm, my discussion is based on the assumption that such a distinction is not physical, and that in reality all systems have an underlying quantum behaviour that may, in some cases, be to some extent suppressed. The process of suppression is sometimes amenable to a description by Ehrenfest’ theorem and by taking trivial limits when $\hbar\rightarrow 0$, but in most cases of interest in the dynamics of neural network, it is not.
@@
One direction of research nowadays is to find some classical underpinnings for quantum mechanics. This research is an important part of the so called ”foundational” research on quantum mechanics. While the common approach is to lift or extend the observables of a system from the uniquely defined classical real numbers to matrix valued observables that encode several possible outcomes without the possibility of predicting which one will occur from the data about the system itself, some quantum foundational research tries to find a classical (therefore well determined and unique) way of characterising the system based on some form of hidden classical variables. My approach is the exact opposite of this path. Instead of searching for some underpinning classical structure to quantum mechanics, my view is that even systems that appear as classical, may still have hidden quantum properties, due, for example to the quantum nature of the transformation maps between the objects of the specific category considered. In particular, systems assumed to be classical, like a neural network, despite being simulated as classical objects in conventional computers, when physically realised, may exhibit quantum phenomena as quasi-classical systems. In particular, the non-linear dynamics of neural networks may in fact amplify quantum behaviour.
@@
There are several macroscopic systems that are generally assumed to be classical, that by the way in which they work, have structures that give them some quantum properties, as far as the definitions above are concerned. One of these systems is the neural network. It is important to mention that there have been essays in which one speculated about the possibility of generating quantum computing processes in big enough biological neural networks by focusing on some detailed mechanism in some molecular substructure of their neural network [1]. I think such an approach is not feasible on a large scale (although one has discovered entanglement of spins in some animal brains [2,3] and substantial progress is being done in understanding macroscopic quantum phenomena from this point of view). Those methods are extremely susceptible to decoherence. However, one may ask what if the large structure of a network has properties that at least in part reflect the fundamental axioms of quantum mechanics?
In this case, the entanglement between spins in animal brains would not be occurring, as described until now, in a substance assimilated to a heat bath, but in an environment adapted to amplify and preserve quantum superpositions by means of the non-linear dynamics resulting from the neural properties it possesses. In this case it would be less surprising to see biological brains taking advantage of quantum superpositions and act like a quantum annealer.
There have also been essays in constructing so called quantum neural networks [4]. Those networks have the links represented by quantum processes implemented by unitary transformations, representing a quantum neural network in the form of a large quantum circuit. Those methods are promising and it would be amazing if they will be achieved. However, there is still the problem of decoherence that will require such networks to be extremely robust, unfortunately a property still beyond today’s reach.
It has however been noticed that classical neural networks do a great job at solving many body quantum problems with strong coupling [5], that are fundamentally quantum and expected to be properly solved on quantum computers. One role neural networks play in solving strongly correlated quantum systems is through simplifying the wavefunction by providing us with a more compact representation thereof. Another role played by neural networks is to improve the understanding of quantum dynamics [12-16]. What are the properties of a classical neural network that would make it suitable for solving such complex quantum problems? While it is true that classical neural networks seem to be very suitable for simulating quantum processes, I wish to ask in this article whether this is due to the fact that the rules of functioning of a neural network that appears classical from the perspective of its objects (namely neurons) has in fact hidden quantum properties at the level of the morphisms (maps) between those objects. Also, it would be interesting to see how the non-linear behaviour of the quasi-classical neural dynamics can affect underlying quantum degrees of freedom.
@@
Indeed, various quantum properties appear to be general properties not only of physical particles, but also of a series of mathematical structures that have a broad applicability. Non-separability, as encountered in the definition of entanglement, is also found, under the guise of non-cartesianity in the construction of fibre bundles, in which basically the bundle measures the global failure of maintaining the local cartesianity of a certain pairing. A non-trivial fibre bundle is defined to characterise precisely such a failure of the cartesian product between the fibre and the base to expand globally. This is just a mathematical example that could make the reasoning I try to follow here clearer: quantum properties are not bound to particles, but can also characterise various rules by which dynamical systems function. Such quantum properties make them function differently compared to their classical counterparts. Of course, this doesn’t mean a classical fibre bundle is quantum. We need additional structure for this. We need some type of quantum polarisation (like in geometric quantisation) and some form of half form correction, both general properties of quantisation. However, there exists a deep similarity between the category of cobordisms whose objects are (n-1)-dimensional manifolds and its morphisms are n-dimensional cobordisms, and the catgeory of Hilb whose objects are Hilbert spaces used to describe states and the morphisms are linear operators describing ”processes” [32]. With this relation, the cartesian non-separability property that leads to entanglement in quantum mechanics has a similar role in other situations that involve non-separability, as is the case in fibre bundles.
From a practical perspective, it is important to know that the same type of entanglement exists in a category related by a functor to the initial one, and it would be possible to harness it as a quantum resource and use it in circumstances that are apparently remote from the initial category. In particular, an apparently classical neural network may extract its entanglement from an associated category and behave like a quantum system when its dynamics is analysed. A wavefunction, or a probability amplitude can be associated not only to the objects of the category, but also to the maps between those objects, and hence superposition and even entanglement described at the level of the wavefunction associated to the quantum maps may prove to be a new quantum resource that could be harnessed by physical neural networks. This would lead to the possibility of having quantum computation like effects occurring even in the most noisy and hot environments, like the ones existing in animal brains. A physical feature emerging from the quantum properties above is the inseparability of state spaces of quantum states [8]. This obstruction to cartesian pairing has been traditionally called entanglement, and it is advisable to keep that notion even when applied to more abstract mathematical objects [9], [32].
@@
Entanglement is considered a characterising feature of quantum mechanics. It is a clearly non-classical phenomenon capable of offering correlations above any that could be obtained classically [10]. Entanglement is based on the non-separability of state spaces of systems that are connected to form an over-arching super-system. The non-separability relies on two concepts. First, if we take two systems, each having their state spaces given, and combine them one obtains a paired state space with a dimension larger than the sum of the dimensions of the two subspaces. The dimension will grow as the product of the dimensions of the two subspaces, leading to a pairing that involves significant global information that cannot be mapped or retrieved by local measurements on any of the two subsystems separately. The fact that pairing is not always cartesian is a feature of quantum mechanics that has correspondences in many algebraic structures. One can link this with the commutation relations that become non-abelian for incompatible observables, and finally with a probabilistic interpretation that allows, by means of statistics, to determine interference patterns linked to the global structure of the manifold on which the phenomena occur. In all cases, the two main aspects of the ”quantumness” of a system are the linear product structure, and its ability to provide access to non-local features by means of a probabilistic, wave-function based interpretation. It is worth noting that the entangled states are, as any quantum state, maximal states of knowledge. They are being defined to describe global properties of the combined system in the case in which separated independent descriptions of each of the subsystems that are combined are unachievable and don’t even exist for the specified observables in an unambiguous sense.
@@
In this article I show that a (quasi)-classical neural network can indeed be represented by a non-cartesian monodromic category and that there exists a functor that links it to the $Hilb$ category. By doing so, a normal (quasi)-classical neural network will show properties of quantum entanglement at a large scale in the same way in which other quantum systems show [37]. However, being highly decoherent at the level of the constituent objects, it will be up to the maps to act as a resource for quantum entanglement. This means that the quantum nature emerges from the rules we impose on how the network should work, not from the actual quantum nature of its constituents, but finally, there should be no difference between a ”natural” quantum system made out of fundamental constituents that are ”quantum”, and one that is forced to function by rules (maps or morphisms of the category) that result in an overall quantum behaviour of the network. The parameters of the network can be encoded in a Hilbert Kernel, where the respective parameters play the role of linear superpositions of the learned results, similar to the quantum superpositions allowed by the complex phases of a wavefunction. Therefore, there appears to be a strong analogy (if not more, a functorial connection) between entangled quantum states and neural networks and between the category of Hilbert spaces and the supposed category of neural networks.
Indeed it may be possible to experimentally test the morphisms origin of some quantum behaviour of a neural network by looking at a physical implementation of a neural network that appears to be classical from the perspective of the objects involved, but may gain quantum properties due to its inherent morphisms. Such physical neural networks have been constructed [17]. Until such an experiment can be performed, I can offer a numerical simulation showing that indeed this is the case, namely that provided with proper input, a physical, classical neural network can entangle the particles that it receives as input.
@@
A (quasi)-classical neural network has certain properties that could be thought of as quantum to some extent. While we do not propagate a quantum state through a series of quantum gates forming a hard to maintain network, we instead use a set of quasi-classical neurons that, acting together, would be able to produce a state that could be classified as quantum for all practical purposes.